Emma is y years old.
Her sister is 4 years younger than Emma.
The product of their ages is 1152.
Show that y^2-4y-1152=0

Answer by Alan3354(69443)

Emma is y years old.
Her sister is 4 years younger than Emma.
The product of their ages is 1152.
Show that y^2-4y-1152=0
------------
y*(y-4) = 1152
y^2 - 4y - 1152 = 0